OpenStudy (anonymous):

How many solutions does the equation 5m – 5m – 12 = 14 – 2 have? (1 point) None One Two Infinite

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Infinity?

OpenStudy (skullpatrol):

Yes, infinitely many.

OpenStudy (skullpatrol):

Do you know why?

OpenStudy (anonymous):

i think so

OpenStudy (skullpatrol):

Explain please :)

OpenStudy (skullpatrol):

5m – 5m – 12 = 14 – 2 (5 - 5)m - 12 = -12 0m = -12 + 12 0m = 0 Zero times any number is zero.